# GRAph-Structured Pyramidal Whole Slide Image Representation

### v1

@@ 4.2 Training and Inference
To tackle the data imbalance problem, for all models in the study, we deployed a weighted cross entropy loss. A learning rate of $0.001$ and a weight decay of $0.01$ for Adam optimizer have been adopted, and in case competing models were not converging, learning rate of $0.0001$ resolved the problem. Models were trained for 100 epochs and 10 epochs for the Ovarian and the Bladder dataset, respectively. Specific to GRASP, the first two layers are of size 256 and the last layer output is of size 128. For all training and testing, the GPU hardware used was either a GeForce GTX 3090 Ti-24 GB (Nvidia) or Quadro RTX 5000-16 GB (Nvidia) based on availability. Deep Graph Library (DGL), PyTorch, NumPy, SciPy, PyGeometric, and Scikit-Learn libraries have been used to perform the experiments.
4.3 Comparisons with State-of-the-Art
To compare with state-of-the-art approaches, we repeated these experiments with the same cross-validation folds and random seeds to have a fair comparison; the choice of ten random seeds is to capture statistical significance and reliability. For evaluating the models, we adopt Balanced Accuracy and F1 Score since these metrics show how reliable a model performs on imbalanced data, and more importantly on clinical applications. We compare our proposed model, GRASP, with models using different approaches to have a broad spectrum of evaluation. These models include Ab-MIL [15], Trans-MIL [25], CLAM-SB [18], and CLAM-MB [18] from the attention/transformer-based family; ZoomMIL [26] from multi-magnification approaches since it has a hierarchical structure and is compatible with our patch extraction paradigm; and PatchGCN: latent & spatial [5] and DGCN: latent & spatial [29] from graph-based learning approaches.
4.4 Subtype Prediction
Table 1 shows the comparison between our model and state-of-the-art methods, where GRASP outperforms all the competing methods by a margin of $2.6\%-10.7\%$ Balanced Accuracy on the Ovarian dataset and $0.4\%-10.0\%$ on the Bladder dataset. It is worth mentioning that ZoomMIL is the closest-performing model to GRASP, although ZoomMIL has 7 times more parameters than GRASP. PatchGCN and DGCN are not performing comparably to GRASP, even though they are using spatial information that GRASP does not. This implies that a multi-magnification graph structure can potentially show more capability compared to other state-of-the-art approaches in terms of representing gigapixel WSIs. Moreover, GRASP and Ab-MIL have the lowest inference time, while GRASP is processing information in three magnifications and Ab-MIL is a single magnification model. Inference times (per slide) have been calculated on the same machine (RTX5000) for all models.
Figure 3: The histogram of consultations conducted by GRASP with different magnifications. First, this shows GRASP is actively dynamic in terms of capturing information from different magnifications benefiting from its multi-magnification structure. Second, information is distributed differently over magnifications depending on the subtype and slide, and there is no optimal magnification for a subtype. For example, in the Bladder dataset, ‘$(5x\&10x\&20x)$’ shows that the model needed to consult with all three magnifications for $19.3\%$ and $39.4\%$ of slides for MicroP and UCC, respectively; ‘$(5x)$’ shows that the model has mostly focused on only $5x$ magnification for $43.2\%$ and $5.6\%$ of slides for MicroP and UCC, respectively. This behavior is similar to pathologists, where they can diagnose massive MicroP tumors with lower magnifications, while they need to consult with higher magnifications for confirming a minuscule mass of MicroP tumors. On the other hand, UCC is hard to diagnose at lower magnifications and requires careful examination with different magnifications due to its morphological complexity, which fits the model behavior in proclivity to highlight more than one magnification for the majority of cases.
Although [7] has used attention score distribution to show their model is reliable across different magnifications, we want to step further and adopt a similar logic as first introduced in [24] to define the concept of energy of gradients for graph nodes (More details in Supplementary Materials). Therefore, for the first time in the field, we show that an AI model such as GRASP can learn the concept of magnification and behave according to the subtype and slide characteristics. To this end, we formulate an experiment to obtain a sense of each magnification’s influence on the model which leads to Figure 3. The main takeaway of this experiment is that depending on the subtype, the distribution of referenced magnifications by GRASP is different. From a pathological point of view, this finding fits our knowledge of the biological properties of each subtype. As an example, we conducted a case study on the bladder dataset, where the micropapillary subtype is known to be diagnosed generally in lower magnification owing to its morphological properties and the structure of micropapillary tumors, whereas UCC needs to be examined in higher magnifications due to its cell- and texture-dependent structure.
@@ 4.5 Ablation Study
Here, we design two experiments to evaluate our proposed model. Firstly, a Monte Carlo test on graph size, i.e., the number of nodes. Secondly, analyzing model performance on individual or pairs of magnifications.
@@ 4.5.1 Monte Carlo Test
In this experiment, we take a graph of size $12003m=$ 1200, and randomly drop a set of its nodes, along with their multi-magnification correspondences, to build a new graph with a smaller size as described in Algorithm 1. For example, when $count=10$, we randomly drop 10 triplets of $(h_{index},h_{index}^{\prime},h_{index}^{\prime\prime})$ from the graph to create a new graph $Q_{r}$ of the size $1170$. Similarly, when $count=200$, we randomly drop 200 triplets of $(h_{index},h_{index}^{\prime},h_{index}^{\prime\prime})$ from the graph to create a new graph $Q_{r}$ of the size $600$. This is an aggressive way to create statistically independent graphs of smaller sizes. To capture statistical variance, we repeat the experiment $10$ times to create $40$ graphs with different sizes and report the model performance in Figure 4. To accomplish this, the same 3-fold cross-validation sets and $10$ random seeds have been used for all repetitions. Taking $10$ times repetitions of $40$ different graph sizes into account, we performed $12,000$ independent training and inference experiments. See Monte Carlo Test in Supplementary Materials.
@@
4:$D\leftarrow\{\emptyset\}$triangleright Node indices to be dropped
5:for 10 times do
@@
7:         for $count$ in $[10:step:390]$ do
8:              $D\leftarrow$ RANDOM$([1,...,m],count)$
9:              $Q_{r,count}\leftarrow G_{r}$
@@
11:                  $Q_{r,count}\leftarrow$ DROP$(Q_{r,count},h_{index})$
12:                  $Q_{r,count}\leftarrow$ DROP$(Q_{r,count},h_{index}^{\prime})$
13:                  $Q_{r,count}\leftarrow$ DROP$(Q_{r,count},h_{index}^{\prime\prime})$
@@
19:Take $Q_{r,count}$ for Cross-Validation and Inference
report Balanced Accuracy & Standard Deviation
Figure 4: Monte Carlo experiment on the graph size.
4.5.2 Magnification Test
Table 2: Average Performance on 3-folds and 10 random seeds
@@ 4.5.2 Magnification Test
To confirm that the idea of multi-magnification is valid and that multi-magnification is the cause for the model’s performance, we design 6 different experiments (repeated on 10 random seeds and 3 folds) on the Bladder dataset as our empirical evidence. These include evaluating the same model on only $\textbf{M}_{1}$, $\textbf{M}_{2}$, and $\textbf{M}_{3}$ fully connected graphs and on pairs of $\textbf{M}_{1}\&\textbf{M}_{2}$, $\textbf{M}_{1}\&\textbf{M}_{3}$, and $\textbf{M}_{2}\&\textbf{M}_{3}$. The results in Table 2 show that GRASP is superior to all other methods. One possible explanation is that for those single and paired graphs, three layers of GCNs most likely cause the aforementioned over-smoothing problem, which shows that GRASP can effectively capture the information contained in different magnifications and boost its performance.
5 Conclusion
In this work, we developed GRASP, a graph-structured multi-magnification framework for processing gigapixel WSIs. GRASP is a dynamic model that can learn multi-magnification interactions in the data based on the idea of capturing both the inter- and intra-magnification information. This relies on the theoretical property of the model, where it benefits from inter-magnification convergence to pool the nodes rather than conventional pooling layers. GRASP, which has comparably fewer parameters than other state-of-the-art multi-magnification models in the field, outperforms all of the competing models in terms of Balanced Accuracy over two complex cancer datasets. For the first time in the field, confirmed by two expert genitourinary pathologists, we showed that our model is dynamic in finding and consulting with different magnifications for subtyping two challenging cancers. We also evaluated the model decision-making to show that the model is learning semantics by highlighting tumorous regions in patches. Furthermore, we not only run extensive experiments to show the model’s reliability and stabilization in terms of its different hyperparameters, but we also provide the theoretical foundation of our work to shed light on the dynamics of GRASP in interacting with different nodes and magnifications in the graph. To conclude, we hope that the strong characteristics of GRASP and its straightforward structure, along with the theoretical basis, will encourage the modeling of structure-based design in the field of digital pathology for WSI representation.
As part of potential future directions, GRASP’s performance on other clinical tasks such as survival analysis and outlier detection, and its potential to incorporate other biological modalities, especially those that are represented as graphs, into GRASP’s graph structure can be investigated.

### v2

@@ 7.2 Training and Inference
To tackle the data imbalance problem, for all models in the study, we deployed a weighted cross entropy loss. A learning rate of $0.001$ and a weight decay of $0.01$ for Adam optimizer have been adopted, and in case competing models were not converging, learning rate of $0.0001$ resolved the problem. Models were trained for 100, 50, and 10 epochs for the Ovarian, ESCA, and Bladder datasets, respectively. Specific to GRASP, the first two layers are of size 256 and the last layer output is of size 128. For all training and testing, the GPU hardware used was either a GeForce GTX 3090 Ti-24 GB (Nvidia), Quadro RTX 5000-16 GB (Nvidia), RTX 6000-48 GB (Nvidia) based on availability. Deep Graph Library (DGL), PyTorch, NumPy, SciPy, PyGeometric, and Scikit-Learn libraries have been used to perform the experiments.
7.3 Theoretical Analysis
Here, we prove Theorem 1 for any $h_{i}^{(3)}$ and $h_{j}^{(3)}$, and conclude the case for $h_{i}^{\prime(3)}$s and $h_{i}^{\prime\prime(3)}$s similarly. To start with, we demonstrate Lemma 1.
Lemma 1.
For any given vectors $x$ and $y$, and having .$\left\|.\right\|_{2}$ as the $L_{2}$ norm, the following inequality holds,
$\centering\left\|\alpha(x)-\alpha(y)\right\|_{2}\leq\left\|x-y\right\|_{2},\@add@centering$ (6)
where $\alpha(.)=\operatorname{ReLU}(.)$
Proof.
Let’s reformulate $\operatorname{ReLU}(x)$ as $\frac{x+\left|x\right|}{2}$ where operator $|x|$ is the element-wise absolute value of the vector $x$. Thus,
$\left\|\alpha(x)-\alpha(y)\right\|_{2}=\left\|\frac{x+|x|}{2}-\frac{y+|y|}{2}\right\|_{2}$
$=\left\|\frac{x-y}{2}+\frac{|x|-|y|}{2}\right\|_{2}$
$\leq\left\|\frac{x-y}{2}\right\|_{2}+\left\|\frac{|x|-|y|}{2}\right\|_{2}$ (7)
using the reverse triangle inequality, $\left\|\frac{|x|-|y|}{2}\right\|_{2}\leq\left\|\frac{x-y}{2}\right\|_{2}$, which gives result to
$\left\|\alpha(x)-\alpha(y)\right\|_{2}\leq\left\|\frac{x-y}{2}\right\|_{2}+\left\|\frac{x-y}{2}\right\|_{2}$
$=\left\|x-y\right\|_{2}$ (8)
Figure 5: The structure of our hierarchical graph and the relationship between two given nodes $h_{i}$ and $h_{k}$ within and across different magnifications.
Theorem 1:
Proof.
Recalling the main GCN formula for any $l$ in 4, $\mathcal{N}(i)$ is the neighborhood size, and taking self-loops (see Note 1) into account, $\forall i\in[1,...,m]$, $\left|\mathcal{N}(i)\right|=m+1$. With the graph being symmetric, we deduce that $\forall j\in[1,...,m]$, $\left|\mathcal{N}(j)\right|=m+1$. These result in
$c_{ji}=\sqrt{\left|\mathcal{N}(j)\right|\left|\mathcal{N}(i)\right|}=m+1$
hence, Eq. 4 is simplified as follows,
$h_{i}^{(l+1)}=\alpha(b^{(l)}+\frac{1}{m+1}\sum_{j\in\mathcal{N}(i)}h_{j}^{(l)}W^{(l)}).$ (9)
Now, for any $i$, we can partition the set of all nodes in $\mathcal{N}(i)$ into two partitions $\{h_{1}^{(l)},...,h_{m}^{(l)}\}$ and $\{h_{i}^{\prime(l)}\}$ based on the relationship between nodes as shown in Figure 5. Therefore, $\sum_{j\in\mathcal{N}(i)}h_{j}^{(l)}$ can be rewritten as,
$\sum_{j\in\mathcal{N}(i)}h_{j}^{(l)}=\left(\sum_{j\in[1,...,m]}h_{j}^{(l)}\right)+h_{i}^{\prime(l)}.$ (10)
The first term in Eq. 10 is common among all nodes at a given magnification, so we call it $\mathcal{H}^{(l)}$ leading to Eq. 11,
$\sum_{j\in\mathcal{N}(i)}h_{j}^{(l)}=\mathcal{H}^{(l)}+h_{i}^{\prime(l)}$ (11)
as a result, we combine Eq. 9 with Eq. 11 which yields
$h_{i}^{(l+1)}=\alpha(b^{(l)}+\frac{1}{m+1}\left(\mathcal{H}^{(l)}+h_{i}^{\prime(l)}\right)W^{(l)})$ (12)
and similarly for any $j\neq i$ as well,
$h_{j}^{(l+1)}=\alpha(b^{(l)}+\frac{1}{m+1}\left(\mathcal{H}^{(l)}+h_{j}^{\prime(l)}\right)W^{(l)}).$ (13)
By using Lemma 1 and with combination with Eq. 12 and 13,
$\left\|h_{i}^{(l+1)}-h_{j}^{(l+1)}\right\|_{2}$
$\leq\left\|\frac{1}{m+1}h_{i}^{\prime(l)}W^{(l)}-\frac{1}{m+1}h_{j}^{\prime(l)}W^{(l)}\right\|_{2}$
$=\frac{1}{m+1}\left\|\left(h_{i}^{\prime(l)}-h_{j}^{\prime(l)}\right)W^{(l)}\right\|_{2}$
$\leq\frac{1}{m+1}\left\|h_{i}^{\prime(l)}-h_{j}^{\prime(l)}\right\|_{2}\left\|W^{(l)}\right\|_{2}.$ (14)
Therefore, we reach the inequality below,
$\left\|h_{i}^{(l+1)}-h_{j}^{(l+1)}\right\|_{2}\leq\frac{1}{m+1}\left\|h_{i}^{\prime(l)}-h_{j}^{\prime(l)}\right\|_{2}\left\|W^{(l)}\right\|_{2}.\$ (15)
Now, by going recursively over $l=0,1,2$, we have
$\left\|h_{i}^{(3)}-h_{j}^{(3)}\right\|_{2}\leq$
$(\frac{1}{m+1})^{3}\left\|h_{i}^{\prime(0)}-h_{j}^{\prime(0)}\right\|_{2}\left\|W^{(2)}\right\|_{2}\left\|W^{(1)}\right\|_{2}\left\|W^{(0)}\right\|_{2}.$ (16)
Since $\left\|W^{(2)}\right\|_{2}$, $\left\|W^{(1)}\right\|_{2}$, and $\left\|W^{(0)}\right\|_{2}$ are $L_{2}-bounded$ based on our assumption. Also, $\left\|h_{i}^{\prime(0)}-h_{j}^{\prime(0)}\right\|_{2}$ is an $L_{2}-bounded$ value based on our assumption (as input image data is $L_{2}-bounded$, and also our encoder $\phi$ is a bounded encoder: features are not scattered in an infinite space, rather they are encoded in a finite space). Given these, by approaching $m\rightarrow\infty$ (see remark 2), the right side of the Eq. 7.3 approaches 00. Therefore,
$\lim_{m\to\infty}\left\|h_{i}^{(3)}-h_{j}^{(3)}\right\|_{2}=0$ (17)
similar to this case, it can be proved that
$\lim_{m\to\infty}\left\|h_{i}^{\prime(3)}-h_{j}^{\prime(3)}\right\|_{2}=0$ (18)
$\lim_{m\to\infty}\left\|h_{i}^{\prime\prime(3)}-h_{j}^{\prime\prime(3)}\right\|_{2}=0$ (19)
$\square$
remark 1.
To implement GCNs, self-loops are considered to represent the relationship between each node with itself, and it is also part of the technical implementation of the models. Thus, we consider this fact in our theoretical discussion.
remark 2.
Empirically, reaching sufficiently large $m$ can guarantee the convergence. For example, $m=10$ can affect the upper bound in Eq. 7.3 with an order of $\frac{1}{10^{3}}$, while $m=100$ can affect the upper bound with an order of $\frac{1}{10^{6}}$. In our experiments, $m=400$ has been adopted that guarantees the convergence with an order of $\frac{1}{64\times 10^{6}}$. Therefore, the larger $m$, the tighter together node features at the last GCN layer.
Corollary 2.
Description: given Eq. 17, every two arbitrary nodes $h_{i}^{(3)}$ and $h_{j}^{(3)}$ in one level of magnification are converging to each other. This means that all nodes are converging to the same value, which we name $h^{*}$. Thus, $\forall i\in[1,...,m]$, $\lim_{m\to\infty}\left\|h_{i}^{(3)}-h^{*}\right\|_{2}=0$ or equivalently $h_{i}^{(3)}\longrightarrow h^{*}$. Using the same logic as above, one can conclude $h_{i}^{\prime(3)}\longrightarrow h^{\prime*}$ and $h_{i}^{\prime\prime(3)}\longrightarrow h^{\prime\prime*}$. Since each of $h^{*}$, $h^{\prime*}$, and $h^{\prime\prime*}$ are a function of $m$, increasing $m$ would result in them being a better estimation/representation for the intra-magnification information.
7.4 Empirical Proof
In addition to the theoretical analysis in Section 7.3, we empirically demonstrate the intra-magnification convergence in Figure 6. In this experiment, we plot the mean squared error between all nodes in a magnification and the corresponding convergence node. As shown, the mean squared error at the third layer ($\ell=3$) is nearly zero, providing empirical evidence for the convergence of the nodes, i.e., the nodes being pooled without the need for a pooling layer.
Figure 6: Empirical proof for intra-magnification convergence at $\ell=3$.
@@ 7.5.1 Monte Carlo Test
As described in Algorithm 1, this test requires $m=400$. Therefore, we removed four slides that did not have 400 non-overlapping patches in $20x$ from the Bladder dataset and ran this test. Because of removing these four slides from the dataset, the result of Figure 4 in the paper is not comparable with any of the tables in the paper. In this test, a subset of the nodes is randomly dropped from each magnification layer (corresponding nodes in each magnification are removed), to create independent graphs. Then, training and inference happens and the results are reported at the end.
@@
4:$D\leftarrow\{\}$triangleright Node indices to be dropped
5:for $\text{iter}\leftarrow 1$ to $10$ do
@@
7:         for $\text{count}\leftarrow 10$ to $390$ step $step$ do
8:              $D\leftarrow\text{RANDOM}([1,\dots,m],\text{count})$
9:              $Q_{r,\text{count}}\leftarrow G_{r}$
@@
11:                  $Q_{r,\text{count}}\leftarrow\text{DROP}(Q_{r,\text{count}},h_{\text{index}})$
12:                  $Q_{r,\text{count}}\leftarrow\text{DROP}(Q_{r,\text{count}},h_{\text{index}}^{\prime})$
13:                  $Q_{r,\text{count}}\leftarrow\text{DROP}(Q_{r,\text{count}},h_{\text{index}}^{\prime\prime})$
@@
19:Use $Q_{r,\text{count}}$ for cross-validation and inference
20:Report: Balanced Accuracy and Standard Deviation
Figure 7: Monte Carlo experiments on the graph size. As the number of nodes increases, the uncertainty decreases and the model stabilizes.
Table 4: Average Performance on 3-folds and 10 random seeds based on KimiaNet’s features.
@@
In this experiment, we take a graph of size $12003m=$ 1200, and randomly drop a set of its nodes, along with their multi-magnification correspondences, to build a new graph with a smaller size. For example, when $count=10$, we randomly drop 10 triplets of $(h_{index},h_{index}^{\prime},h_{index}^{\prime\prime})$ from the graph to create a new graph $Q_{r}$ of the size $1170$. Similarly, when $count=200$, we randomly drop 200 triplets of $(h_{index},h_{index}^{\prime},h_{index}^{\prime\prime})$ from the graph to create a new graph $Q_{r}$ of the size $600$. This is an aggressive way to create statistically independent graphs of smaller sizes. To capture statistical variance, we repeat the experiment $10$ times to create $40$ graphs with different sizes and report the model performance in Figure 7. To accomplish this, the same 3-fold cross-validation sets and $10$ random seeds have been used for all repetitions. Taking $10$ times repetitions of $40$ different graph sizes into account, we performed $12,000$ independent training and inference experiments. Algorithm 1 demonstrates this experiment.
7.5.2 Magnification Test
@@
To confirm that the idea of multi-magnification is valid and that multi-magnification is the cause for the model’s performance, we design 6 different experiments (repeated on 10 random seeds and 3 folds) on the Bladder dataset, with KimiaNet as the backbone, as our empirical evidence. These include evaluating the same model on only $\textbf{M}_{1}$, $\textbf{M}_{2}$, and $\textbf{M}_{3}$ fully connected graphs and on pairs of $\textbf{M}_{1}\&\textbf{M}_{2}$, $\textbf{M}_{1}\&\textbf{M}_{3}$, and $\textbf{M}_{2}\&\textbf{M}_{3}$. The results in Table 4 show that GRASP is superior to all other methods. One possible explanation is that for those single and paired graphs, three layers of GCNs most likely cause the aforementioned over-smoothing problem, which shows that GRASP can effectively capture the information contained in different magnifications and boost its performance.
7.5.3 Graph Convolutions
To study the impact of different graph convolutions on the performance of GRASP, we designed this experiment where we replaced the GCN layers in the original layers, with a newer version of graph convolutions. As seen in Figure 8, GAT and SAGEConv improve the performance of the primary GCN-based GRASP in terms of Balanced Accuracy over different batch sizes.
Figure 8: Ablation on the effect of different GNN structures benchmarked on 3 folds and 10 random seeds with different batch sizes. GNNs were taken from Deep Graph Library.
Table 5: The average performance on 3 folds and 10 random seeds, based on Swin’s features and the setting where all the patches were extracted from each WSI.
Model Bladder: Two Subtypes
Balanced Acc. F1 Score AUC
ZoomMIL $0.879\pm 0.065$ $0.872\pm 0.060$ $0.951\pm 0.031$
HiGT $0.720\pm 0.049$ $0.658\pm 0.042$ $0.819\pm 0.066$
H2MIL $0.877\pm 0.050$ $0.871\pm 0.035$ $0.966\pm 0.022$
GRASP (GCN) $\mathbb{0.883\pm 0.069}$ $\mathbb{0.879\pm 0.065}$ $\mathbb{0.953\pm 0.031}$
GRASP (GAT) $\mathbb{0.917\pm 0.013}$ $\mathbb{0.907\pm 0.017}$ $\mathbb{0.978\pm 0.007}$
GRASP (SAGEConv) $\mathbb{0.936\pm 0.023}$ $\mathbb{0.932\pm 0.015}$ $\mathbb{0.988\pm 0.008}$
Figure 9: Our model’s performance as the number of GCN layers increases. As can be seen, by increasing the number of layers from three to four a relatively large drop happens, showing that the model is being over-smoothed. However, the model is recovering this loss at $five$ layers or more yet with a relatively larger standard deviation compared to three layers.
7.5.4 Patch Number
To compare the models’ performance when all the patches from each slide are extracted, we designed this experiment and benchmarked the baseline against different variations of GRASP. GCN-based GRASP is superior to other multi-magnification methods such as ZoomMIL, H2MIL, and HiGT. Compared to Table 5, the increased number of patches helps GRASP to improve, yet it degrades the performance of the other three models. However, single Magnification methods are more competitive with a higher density of patches. With this in mind, we followed the previous ablation study 7.5.3 and compared GRASP’s performance with newer graph convolutions. Consequently, GRASP with SAGEConv outperforms all other models. This further emphasizes the flexibility of the graph structure that can easily employ different graph convolutions, which we belive is a unique advantage of GRASP.
7.5.5 Graph Depth
We experimented with GRASP with different numbers of GCN layers as shown in Figure 9. Firstly, three layers of GCNs show the same performance as two layers yet with lower standard deviation. Secondly, four layers of GCNs show a sudden drop in performance and increase in standard deviation, which can be attributed to over smoothing problem Cai & Wang (2020); Wu et al. (2023). In addition, with more than $five$ layers of GCNs, the network can recover the performance but with a slightly higher standard deviation and, clearly, an increased number of parameters. This shows that our original architecture of three layers is the best choice in the trade-off of average accuracy and the model’s reliability. Based on the discussion in Wu et al. (2023), we expect the same phenomenon for different graph convolution types to happen.
7.6 Graph-Based Visualization
Let’s call the output of the classifier, $S$, where the logit for correctly classified slides/graphs is $S_{c}$. To visualize the importance of magnifications, we compute the magnitude of the gradient of a graph with respect to its node features at $l=0$ ($h_{i}^{(0)}$; for the sake of brevity, we drop the superscript $(0)$ form $h_{i}^{(0)}$ and show it as $h_{i}$), which is $\left|\frac{\partial S_{c}}{\partial h_{i}}\right|$ for the magnification $\textbf{M}_{1}$. $\left|\frac{\partial S_{c}}{\partial h_{i}}\right|$ is a vector of size $1024\times 1$, and we have $m$ nodes giving result to $m$ such vectors. Likewise, we can define $\left|\frac{\partial S_{c}}{\partial h_{i}^{\prime}}\right|$ and $\left|\frac{\partial S_{c}}{\partial h_{i}^{\prime\prime}}\right|$ for $\textbf{M}_{2}$ and $\textbf{M}_{3}$, respectively. Arranging these absolute gradients for each magnification in a matrix of size $m\times 1024$ as follows,
$\textit{Heatmap}_{\mathbf{M}_{1}}=\left[\begin{array}[]{c}\left|\frac{\partial S_{c}}{\partial h_{1}}\right|\\ \vdots\\ \left|\frac{\partial S_{c}}{\partial h_{m}}\right|\end{array}\right]$ (20)
$\textit{Heatmap}_{\mathbf{M}_{2}}=\left[\begin{array}[]{c}\left|\frac{\partial S_{c}}{\partial h_{1}^{\prime}}\right|\\ \vdots\\ \left|\frac{\partial S_{c}}{\partial h_{m}^{\prime}}\right|\end{array}\right]$ (21)
$\textit{Heatmap}_{\mathbf{M}_{3}}=\left[\begin{array}[]{c}\left|\frac{\partial S_{c}}{\partial h_{1}^{\prime\prime}}\right|\\ \vdots\\ \left|\frac{\partial S_{c}}{\partial h_{m}^{\prime\prime}}\right|\end{array}\right]$ (22)
As such, putting matrices in 20, 21, and 22 together in a matrix gives us the overall heatmap for the graph, of the size $3m\times 1024$, to compare the influence of each magnification:
$\textit{Heatmap}=\left[\begin{array}[]{c}\textit{Heatmap}_{\mathbf{M}_{1}}\\ \textit{Heatmap}_{\mathbf{M}_{2}}\\ \textit{Heatmap}_{\mathbf{M}_{3}}\end{array}\right]$ (23)
This is the heatmap depicted in Figure 9 as a graph-based heatmap, which shows how model focuses on different magnifications.
Having the gradient for each node in the graph, we develop the concept of energy of gradients to find out which magnification(s) play a more important role in GRASP’s final decision. To do so, we start by defining $\mathcal{E}_{\mathbf{M}_{1}}$ as follows,
$\mathcal{E}_{\mathbf{M}_{1}}=\sum_{i\in[1,...,m]}\left\|\frac{\partial S_{c}}{\partial h_{i}}\right\|_{2}^{2}$ (24)
similarly, for $\mathbf{M}_{2}$ and $\mathbf{M}_{3}$,
$\mathcal{E}_{\mathbf{M}_{2}}=\sum_{i\in[1,...,m]}\left\|\frac{\partial S_{c}}{\partial h_{i}^{\prime}}\right\|_{2}^{2}$ (25)
$\mathcal{E}_{\mathbf{M}_{3}}=\sum_{i\in[1,...,m]}\left\|\frac{\partial S_{c}}{\partial h_{i}^{\prime\prime}}\right\|_{2}^{2}$ (26)
Having these energies, the energy contribution of each magnification is calculated based on their relative share in the whole energy spent in the graph:
${\mathbf{M}_{1}}\textit{'s contribution}=\frac{\mathcal{E}_{\mathbf{M}_{1}}}{\mathcal{E}_{\mathbf{M}_{1}}+\mathcal{E}_{\mathbf{M}_{2}}+\mathcal{E}_{\mathbf{M}_{3}}}$ (27)
${\mathbf{M}_{2}}\textit{'s contribution}=\frac{\mathcal{E}_{\mathbf{M}_{2}}}{\mathcal{E}_{\mathbf{M}_{1}}+\mathcal{E}_{\mathbf{M}_{2}}+\mathcal{E}_{\mathbf{M}_{3}}}$ (28)
${\mathbf{M}_{3}}\textit{'s contribution}=\frac{\mathcal{E}_{\mathbf{M}_{3}}}{\mathcal{E}_{\mathbf{M}_{1}}+\mathcal{E}_{\mathbf{M}_{2}}+\mathcal{E}_{\mathbf{M}_{3}}}$ (29)
Accordingly, the importance of each magnification can be quantified for further investigations. More samples are provides in Figure 3.